# Data-driven Implementations of Various Generalizations of Balanced Truncation

### v2

@@ 2.4 Balanced Truncation (BT) Family
These Gramians solve the following Lyapunov equations:
@@
The Cholesky factorizations of the Gramians are given by
$P=L_{p}L_{p}^{T}\quad\text{and}\quad Q=L_{q}L_{q}^{T}.$
The balanced square-root algorithm (BSA) [33] is among the most numerically stable methods for implementing BT [9]. It begins by computing the singular value decomposition (SVD) of the product $L_{q}^{T}L_{p}$:
@@
The projection matrices $\hat{W}$ and $\hat{V}$ in BSA are constructed as:
@@
The ROM constructed by BT preserves the stability of the original system $G(s)$ as well as its $r$ largest Hankel singular values $\sqrt{\lambda_{i}(PQ)}$. To preserve additional properties—such as stable minimum phase, positive-realness, or bounded-realness—generalized versions of BT have been developed, which primarily differ in their definitions of the Gramians.
In LQG-BT [26], the Gramian-like matrices are computed by solving the following filter and controller Riccati equations:
@@
In $\mathcal{H}_{\infty}$-BT [27], the Gramian-like matrices are computed by solving the following filter and controller Riccati equations with $\gamma>0$:
@@
In PR-BT [24, 34], the Gramians solve the following Riccati equations:
@@
In BR-BT [35, 34], the Gramians solve the following Riccati equations:
@@
In SW-BT [28], the controllability Gramian $P$ remains the same as in BT, while the weighted observability Gramian $Q_{\text{SW}}$ solves the following Lyapunov equation:
@@
In BST [24, 36], the controllability Gramian $P$ remains the same as in BT, while the weighted observability Gramian $Q_{\text{S}}$ solves the following Riccati equation:
@@
By replacing $Q$ with $Q_{\text{S}}$ in BT, a ROM that preserves the minimum-phase property of the original model can be obtained. Unlike SW-BT, BST can handle non-minimum-phase models as well. Both SW-BT and BST minimize the relative error $G^{-1}(s)(G(s)-\hat{G}(s))$ and tend to ensure uniform accuracy across the entire frequency spectrum.
2.5 Quadrature-based Balanced Truncation (QuadBT) [1]
The integrals (12) and (13) can be approximated numerically using quadrature rules as:
@@ 2.5 Quadrature-based Balanced Truncation (QuadBT) [1]
where $\sigma_{i}$ and $\mu_{i}$ are the quadrature nodes, and $w_{p,i}^{2}$ and $w_{q,i}^{2}$ are their corresponding weights. The factorizations $\tilde{P}=\tilde{L}_{p}\tilde{L}_{p}^{*}$ and $\tilde{Q}=\tilde{L}_{q}\tilde{L}_{q}^{*}$ can be expressed as: $\tilde{L}_{p}=\hat{V}\hat{L}_{p}$ and $\tilde{L}_{q}=\hat{W}\hat{L}_{q}$, with $\hat{V}$ and $\hat{W}$ as in (1) and (2), respectively, and
@@
Here, $\hat{L}_{p}$ and $\hat{L}_{q}$ depend only on the quadrature weights. Moreover, the terms $\hat{W}^{*}\hat{V}$, $\hat{W}^{*}A\hat{V}$, $\hat{W}^{*}B$, $C\hat{V}$, and $D$ can be constructed non-intrusively using samples of $G(s)$ via the Loewner framework. In the BSA, $\tilde{L}_{p}$ and $\tilde{L}_{q}$ replace $L_{p}$ and $L_{q}$, leading to the decomposition:
@@
The projection matrices $W_{r}$ and $V_{r}$ are then defined as:
@@
Finally, the ROM in QuadBT is constructed as:
@@
    In PR-BT, the Gramians $P_{\text{PR}}$ and $Q_{\text{PR}}$ are respectively the controllability Gramian and observability Gramian associated with state-space realizations of the spectral factorizations of $G(s)+G^{*}(s)$.
@@
    In BR-BT, the Gramians $P_{\text{BR}}$ and $Q_{\text{BR}}$ are respectively the controllability Gramian and observability Gramian associated with state-space realizations of the spectral factorizations of $I_{m}-G^{*}(s)G(s)$ and $I_{p}-G(s)G^{*}(s)$.
@@
    The QuadBT framework could be extended to PR-BT, BR-BT, and BST if samples were available for the spectral factorizations of $G(s)+G^{*}(s)$, $I_{m}-G^{*}(s)G(s)$, $I_{p}-G(s)G^{*}(s)$, and $G(s)G^{*}(s)$.
@@
2.6 Data-driven ADI-based Balanced Truncation (DD-ADI-BT) [2]
In [37], it is shown that PORK and the ADI method [16] produce identical approximations of Lyapunov equations when the ADI shifts are selected as the mirror images of the interpolation points. Consequently, the ADI-based approximation of $P$ for the shifts $(-\sigma_{1},\dots,-\sigma_{v})$ can be obtained via I-PORK as $P\approx\tilde{P}=\hat{V}Q_{v}^{-1}\hat{V}^{*}$. Similarly, the approximation of $Q$ for the shifts $(-\mu_{1},\dots,-\mu_{w})$ can be computed via O-PORK as $Q\approx\tilde{Q}=\hat{W}P_{w}^{-1}\hat{W}^{*}$. By factorizing $Q_{v}^{-1}=\hat{L}_{p}\hat{L}_{p}^{*}$ and $P_{w}^{-1}=\hat{L}_{q}\hat{L}_{q}^{*}$, and defining $\tilde{L}_{p}=\hat{V}\hat{L}_{p}$ and $\tilde{L}_{q}=\hat{W}\hat{L}_{q}$, the approximations $\tilde{P}$ and $\tilde{Q}$ can be expressed in factorized form as $\tilde{P}=\tilde{L}_{p}\tilde{L}_{p}^{*}$ and $\tilde{Q}=\tilde{L}_{q}\tilde{L}_{q}^{*}$. As noted in [2], replacing $L_{p}$ and $L_{q}$ with $\tilde{L}_{p}$ and $\tilde{L}_{q}$ in BSA yields a non-intrusive implementation of BT. This is because $\hat{L}_{p}$ and $\hat{L}_{q}$ depend only on the interpolation points, and the quantities $\hat{W}^{*}\hat{V}$, $\hat{W}^{*}A\hat{V}$, $\hat{W}^{*}B$, $C\hat{V}$, and $D$ can be computed non-intrusively within the Loewner framework. Thus, the ADI-based BT algorithm can be implemented non-intrusively using transfer function samples in the right half-plane, unlike QuadBT, which requires samples along the imaginary axis.
3 Main Work
In this section, we present several new results on data-driven MOR and controller design, outlined as follows:
  1. 1.
    Recall that the interpolant $\hat{G}(s)=C\hat{V}(sI-S_{v}+\zeta L_{v})^{-1}\zeta+D$ interpolates $G(s)$ at $(\sigma_{1},\cdots,\sigma_{v})$, where $\hat{V}$ is defined in (1) and $\zeta$ is a free parameter. We identify a specific choice of $\zeta$ such that the RADI [4]-based approximation of $P_{\text{LQG}}$ for ADI shifts $(-\sigma_{1},\ldots,-\sigma_{v})$ can be produced by solving a Lyapunov equation involving only the interpolation points $(\sigma_{1},\ldots,\sigma_{v})$ and the product $C\hat{V}$.
  2. 2.
    Similarly, for the interpolant $\hat{G}(s)=\zeta(sI-S_{w}+L_{w}\zeta)^{-1}\hat{W}^{*}B+D$, which interpolates $G(s)$ at $(\mu_{1},\ldots,\mu_{w})$ with $\hat{W}$ defined in (2), we provide a specific $\zeta$ such that the RADI-based approximation of $Q_{\text{LQG}}$ for ADI shifts $(-\mu_{1},\ldots,-\mu_{w})$ can be computed by solving a Lyapunov equation involving only the interpolation data $(\mu_{1},\ldots,\mu_{w})$ and $\hat{W}^{*}B$.
  3. 3.
    Using these approximations of $P_{\text{LQG}}$ and $Q_{\text{LQG}}$, we develop a non-intrusive implementation of LQG-BT, which enables LQG controller design based solely on samples of the plant’s transfer function $G(s)$ in the right-half of the $s$-plane.
  4. 4.
    A similar approach is proposed for the non-intrusive implementation of $\mathcal{H}_{\infty}$-BT, which enables $\mathcal{H}_{\infty}$ controller design based solely on samples of the plant’s transfer function $G(s)$ in the right-half of the $s$-plane.
  5. 5.
    The work in [38] presents linear matrix inequalities (LMIs)- and Riccati-based methods for choosing $\zeta$ in $\hat{G}(s)=C\hat{V}(sI-S_{v}+\zeta L_{v})^{-1}\zeta+D$ that preserve positive-realness, bounded-realness, and minimum-phase properties. However, the formulas in [38] are intrusive since they require access to the state-space matrices $(A,B,C,D)$, and they are also computationally expensive. In contrast, we propose non-intrusive methods to compute $\zeta$ in both $\hat{G}(s)=C\hat{V}(sI-S_{v}+\zeta L_{v})^{-1}\zeta+D$ and $\hat{G}(s)=\zeta(sI-S_{w}+L_{w}\zeta)^{-1}\hat{W}^{*}B+D$ that preserve these system properties. These new formulas require only the transfer function samples and can be derived using Lyapunov and Sylvester equations (or even directly from $G(s)$, as discussed later).
  6. 6.
    We also provide non-intrusive formulas for computing $\zeta$ in the above interpolants that yield the same approximations of $P_{\text{PR}}$, $Q_{\text{PR}}$, $P_{\text{BR}}$, $Q_{\text{BR}}$, $Q_{\text{S}}$ as those produced by RADI. Together with the approximation of $Q_{\text{SW}}$ obtained via O-PORK, we construct non-intrusive implementations of PR-BT, BR-BT, SW-BT, and BST that rely solely on transfer function samples in the right-half of the $s$-plane.
3.1 RADI-based Non-intrusive Implementation of LQG-BT
@@
As demonstrated in [37], the ADI method for computing the low-rank solution of $P$ with shifts $(-\sigma_{1},\dots,-\sigma_{v})$ corresponds to a Petrov-Galerkin approach. This method implicitly interpolates at $(\sigma_{1},\dots,\sigma_{v})$ using the projection matrix $\hat{V}$, while the projection matrix $\hat{W}$ is implicitly chosen to satisfy $\hat{W}^{*}\hat{V}=I$, ensuring that the poles of $\hat{A}$ are placed at $(-\sigma_{1}^{*},\dots,-\sigma_{v}^{*})$. Similarly, [39] shows that the low-rank solution of $P_{\text{LQG}}$ obtained via RADI with shifts $(-\sigma_{1},\dots,-\sigma_{v})$ can be interpreted as a Petrov-Galerkin method. Here, interpolation is implicitly performed at $(\sigma_{1},\dots,\sigma_{v})$ using $\hat{V}$, and $\hat{W}$ is selected such that $\hat{W}^{*}\hat{V}=I$, placing the poles of $\hat{A}-\hat{P}_{\text{LQG}}\hat{C}^{*}\hat{C}$ at $(-\sigma_{1}^{*},\dots,-\sigma_{v}^{*})$. The matrix $\hat{P}_{\text{LQG}}$ solves the projected Riccati equation:
@@
    Given $\hat{A}=\hat{P}_{\text{LQG}}(-S_{v}^{*}+\hat{P}_{\text{LQG}}\hat{C}^{*}\hat{C})\hat{P}_{\text{LQG}}^{-1}$, the eigenvalues of $\hat{A}$ are equal to those of $S_{v}^{*}+\hat{P}_{\text{LQG}}\hat{C}^{*}\hat{C}$. As $S_{v}^{*}$ is Hurwitz, the matrix $\hat{A}-\hat{P}_{\text{LQG}}\hat{C}^{*}\hat{C}$ is Hurwitz with eigenvalues at $(-\sigma_{1}^{*},\cdots,-\sigma_{v}^{*})$.
It is clear from Theorem 3.1 that the RADI-based approximation of $P_{\text{LQG}}$ is given by $P_{\text{LQG}}\approx\hat{V}Q_{v}^{-1}\hat{V}^{*}$.
Note that $Q_{\text{LQG}}$ is the dual of $P_{\text{LQG}}$. For completeness, we present the dual of Theorem 3.1 below, which produces the same approximation of $Q_{\text{LQG}}$ as RADI when the shifts are chosen as $(-\mu_{1},\cdots,-\mu_{w})$.
Theorem 3.2.
@@
The proof is similar to that of Theorem 3.1 and hence omitted for brevity. ∎
Again, it is clear from Theorem 3.2 that the RADI-based approximation of $Q_{\text{LQG}}$ is given by $Q_{\text{LQG}}\approx\hat{W}P_{w}^{-1}\hat{W}^{*}$.
Similar to DD-ADI-BT [2], the RADI-based non-intrusive implementation of LQG-BT can be directly obtained using the results of Theorems 3.1 and 3.2. The pseudo-code for this implementation, named “Data-driven LQG-BT (DD-LQG-BT)”, is provided in Algorithm 1.
Algorithm 1 $D$-LQG-BT
Input: RADI shifts for approximating $P_{\text{LQG}}$: $(-\sigma_{1},\cdots,-\sigma_{v})$; RADI shifts for approximating $Q_{\text{LQG}}$: $(-\mu_{1},\cdots,-\mu_{w})$; Data: $\big(H(\sigma_{1}),\cdots,H(\sigma_{v}),H(\mu_{1}),\cdots,H(\mu_{w})\big)$ and $H^{\prime}(\sigma_{i})$ for $\sigma_{i}=\mu_{j}$; Reduced order: $r$.
[…]
2: Set $S_{v}$, $L_{v}$, $S_{w}$, $L_{w}$ as in (6).
3: Compute $Q_{v}$ and $P_{w}$ by solving the Lyapunov equations (30) and (31).
4: Decompose $Q_{v}^{-1}$ and $P_{w}^{-1}$ as $Q_{v}^{-1}=\hat{L}_{p}\hat{L}_{p}^{*}$ and $P_{w}^{-1}=\hat{L}_{q}\hat{L}_{q}^{*}$.
5: Compute $\hat{V}_{r}$ and $\hat{W}_{r}$ from (26) and (27).
6: Compute $\hat{A}$, $\hat{B}$, and $\hat{C}$ from (28).
@@
$\mathcal{H}_{\infty}$-BT is similar to LQG-BT with the only difference being in the computation of $Q_{v}$ and $P_{w}$. In $\mathcal{H}_{\infty}$-BT, $Q_{v}$ and $P_{w}$ are computed by solving the following Lyapunov equations:
@@
All other steps in Algorithm 1 remain the same.
3.2 RADI-based Non-intrusive Implementation of PR-BT

### v3

@@ 2.4 Balanced Truncation (BT) and Its extensions
These Gramians are solutions to the Lyapunov equations
@@
If the Cholesky factorizations $P=L_{p}L_{p}^{T}$ and $Q=L_{q}L_{q}^{T}$ are computed, the balanced square-root algorithm (BSA) [53] provides a numerically robust way to perform BT [34]. The method starts with the singular value decomposition (SVD) of $L_{q}^{T}L_{p}$:
@@
From this decomposition, the projection matrices are obtained as
@@
The resulting ROM preserves the asymptotic stability of the original model $G(s)$ as well as its $r$ largest Hankel singular values $\sqrt{\lambda_{i}(PQ)}$. To preserve additional properties—such as stable minimum phase, positive-realness, or bounded-realness—generalized versions of BT have been developed, which primarily differ in their definitions of the Gramians.
In LQGBT [26], the Gramian-like matrices are computed as stabilizing solutions to the following filter and controller Riccati equations:
@@
In $\mathcal{H}_{\infty}$BT [33], the Gramian-like matrices are computed as stabilizing solutions to the following filter and controller Riccati equations with $\gamma>0$:
@@
In PRBT [20, 41], the Gramians are computed as stabilizing solutions to the following Riccati equations:
@@
In BRBT [44, 41], the Gramians are computed as stabilizing solutions to the following Riccati equations:
@@
In SWBT [60], the controllability Gramian $P$ remains the same as in BT, while the weighted observability Gramian $Q_{\text{SW}}$ solves the following Lyapunov equation:
@@
In BST [20, 12], the controllability Gramian $P$ remains the same as in BT, while the weighted observability Gramian $Q_{\text{S}}$ is computed as stabilizing solution to the following Riccati equation:
@@
By replacing $Q$ with $Q_{\text{S}}$ in BT, a ROM that preserves the minimum-phase property of the original model can be obtained. Unlike SWBT, BST can handle non-minimum-phase models as well. Both SWBT and BST minimize the relative error
$G^{-1}(s)\Big(G(s)-\hat{G}(s)\Big)$
and tend to ensure uniform accuracy across the entire frequency spectrum.
2.5 Quadrature-based BT (QuadBT) [16]
The frequency-domain integrals in (12) and (13) can be approximated by numerical quadrature as:
@@ 2.5 Quadrature-based BT (QuadBT) [16]
where $\sigma_{i}$ and $\mu_{i}$ denote the quadrature nodes, and $w_{p,i}^{2}$ and $w_{q,i}^{2}$ are the associated weights. These approximations admit Cholesky-like factorizations $\tilde{P}=\tilde{L}_{p}\tilde{L}_{p}^{*}$ and $\tilde{Q}=\tilde{L}_{q}\tilde{L}_{q}^{*}$, which can be written as
$\tilde{L}_{p}=\hat{V}\hat{L}_{p},\qquad\tilde{L}_{q}=\hat{W}\hat{L}_{q},$
with $\hat{V}$ and $\hat{W}$ defined as in (1) and (2), respectively, and where
@@
In BSA, the exact Cholesky factors $L_{p}$ and $L_{q}$ are replaced by $\tilde{L}_{p}$ and $\tilde{L}_{q}$, resulting in the SVD
@@
From this decomposition, the projection matrices are computed as:
@@
Finally, the ROM in QuadBT is obtained as:
@@
Note that $\hat{L}_{p}$ and $\hat{L}_{q}$ depend solely on the quadrature weights. Furthermore, the quantities $\hat{W}^{*}\hat{V}$, $\hat{W}^{*}A\hat{V}$, $\hat{W}^{*}B$, $C\hat{V}$, and $D$ can be constructed in a non-intrusive manner using frequency-domain samples of the transfer function $G(s)$ within the Loewner framework (5). Thus, the ROM in (28) can be constructed non-intrusively, without access to the state-space realization $(A,B,C,D)$.
@@
    In PRBT, the Gramians $P_{\text{PR}}$ and $Q_{\text{PR}}$ are respectively the controllability Gramian and observability Gramian associated with state-space realizations of the spectral factorizations of $G(s)+G^{*}(s)$.
@@
    In BRBT, the Gramians $P_{\text{BR}}$ and $Q_{\text{BR}}$ are respectively the controllability Gramian and observability Gramian associated with state-space realizations of the spectral factorizations of $I_{m}-G^{*}(s)G(s)$ and $I_{p}-G(s)G^{*}(s)$.
@@
    The QuadBT framework could be extended to PRBT, BRBT, and BST if samples were available for the spectral factorizations of $G(s)+G^{*}(s)$, $I_{m}-G^{*}(s)G(s)$, $I_{p}-G(s)G^{*}(s)$, and $G(s)G^{*}(s)$.
@@
2.6 Non-intrusive ADI-based Approximation of BT [61]
In [57], it is demonstrated that the PORK algorithm and the low-rank Cholesky factor ADI (LRCF-ADI) method [8] yield identical low-rank approximations of the solutions to Lyapunov equations, provided the ADI shift parameters are chosen as the mirror images of the interpolation points used in PORK. Specifically, when the ADI shifts are set to $(-\sigma_{1},\dots,-\sigma_{v})$, the resulting approximation of the controllability Gramian $P$ coincides with the I-PORK output and can be written as
$P\approx\tilde{P}=\hat{V}Q_{v}^{-1}\hat{V}^{*}.$
Likewise, for the observability Gramian $Q$, using ADI shifts $(-\mu_{1},\dots,-\mu_{w})$ leads to the O-PORK approximation
$Q\approx\tilde{Q}=\hat{W}P_{w}^{-1}\hat{W}^{*}.$
By computing Cholesky-like factorizations $Q_{v}^{-1}=\hat{L}_{p}\hat{L}_{p}^{*}$ and $P_{w}^{-1}=\hat{L}_{q}\hat{L}_{q}^{*}$, and defining the low-rank factors $\tilde{L}_{p}=\hat{V}\hat{L}_{p}$ and $\tilde{L}_{q}=\hat{W}\hat{L}_{q}$, the approximations admit the factorized representations $\tilde{P}=\tilde{L}_{p}\tilde{L}_{p}^{*}$ and $\tilde{Q}=\tilde{L}_{q}\tilde{L}_{q}^{*}$. Note that, unlike QuadBT, $\hat{L}_{p}$ and $\hat{L}_{q}$ here are not diagonal matrices. As observed in [61], substituting the exact Gramian factors $L_{p},L_{q}$ in BSA with these approximations $\tilde{L}_{p},\tilde{L}_{q}$ yields a non-intrusive ADI-based approximation of BT (NI-ADI-BT). This is because $\hat{L}_{p}$ and $\hat{L}_{q}$ depend solely on the interpolation points, and all required projected system quantities—namely $\hat{W}^{*}\hat{V}$, $\hat{W}^{*}A\hat{V}$, $\hat{W}^{*}B$, $C\hat{V}$, and DD—can be assembled directly from transfer function evaluations using the Loewner framework (5), without access to the original state-space matrices $(A,B,C,D)$. Consequently, NI-ADI-BT can be implemented using transfer function samples in the right half of the $s$-plane, in contrast to QuadBT, which relies on samples along the $j\omega$ axis.
3 Non-intrusive ADI-based Approximations of Various Generalizations of BT
Recall that the interpolant $\hat{G}(s)=C\hat{V}(sI-S_{v}+\zeta L_{v})^{-1}\zeta+D$ interpolates $G(s)$ at $(\sigma_{1},\cdots,\sigma_{v})$, where $\hat{V}$ is defined in (1) and $\zeta$ is a free parameter. Similarly, the interpolant $\hat{G}(s)=\zeta(sI-S_{w}+L_{w}\zeta)^{-1}\hat{W}^{*}B+D$ interpolates $G(s)$ at $(\mu_{1},\ldots,\mu_{w})$, where $\hat{W}$ defined in (2). This section derives specific choices of free parameter $\zeta$ so that rational interpolation produces approximations of various Lyapunov and Riccati equations defining Gramians in various generalizations of BT such that these approximations are identical to ones produced by ADI methods for Lyapunov equations [8] and Riccati equations [3]. Furthermore, when these approximations are inserted into BSA, it is noted that the resulting ROM in various generalizations of BT can be obtained non-intrusively from the samples of $G(s)$ in the right-half of the $s$-plane. The parameter $\zeta$ can be computed by solving specific Lyapunov and Sylvester equations according to the specific generalization of BT.
The following assumptions are made on the $D$ matrix for PRBT, BRBT, SWBT, and BST. For PRBT, $D+D^{T}>0$. For BRBT, $I-DD^{T}>0$ and $I-D^{T}D>0$. For SWBT and BST, $DD^{T}>0$.
3.1 Non-intrusive ADI-based Approximation of LQGBT
@@
As demonstrated in [57], the LRCF-ADI method for computing the low-rank solution of $P$ with shifts $(-\sigma_{1},\dots,-\sigma_{v})$ corresponds to a Petrov-Galerkin approach. This method implicitly interpolates at $(\sigma_{1},\dots,\sigma_{v})$ using the projection matrix $\hat{V}$, while the projection matrix $\hat{W}$ is implicitly chosen to satisfy $\hat{W}^{*}\hat{V}=I$, ensuring that the poles of $\hat{A}$ are placed at $(-\sigma_{1}^{*},\dots,-\sigma_{v}^{*})$. Similarly, [5] shows that the low-rank solution of $P_{\text{LQG}}$ obtained via ADI method for Riccati equations (RADI) with shifts $(-\sigma_{1},\dots,-\sigma_{v})$ can be interpreted as a Petrov-Galerkin method. Here, interpolation is implicitly performed at $(\sigma_{1},\dots,\sigma_{v})$ using $\hat{V}$, and $\hat{W}$ is selected such that $\hat{W}^{*}\hat{V}=I$, placing the poles of $\hat{A}-\hat{P}_{\text{LQG}}\hat{C}^{*}\hat{C}$ at $(-\sigma_{1}^{*},\dots,-\sigma_{v}^{*})$. The matrix $\hat{P}_{\text{LQG}}$ solves the projected Riccati equation:
@@
    Given $\hat{A}=-\hat{P}_{\text{LQG}}S_{v}^{*}\hat{P}_{\text{LQG}}^{-1}+\hat{P}_{\text{LQG}}\hat{C}^{*}\hat{C}$, the eigenvalues of $\hat{A}$ are equal to those of $S_{v}^{*}+\hat{P}_{\text{LQG}}\hat{C}^{*}\hat{C}$. As $S_{v}^{*}$ is Hurwitz, the matrix $\hat{A}-\hat{P}_{\text{LQG}}\hat{C}^{*}\hat{C}$ is Hurwitz with eigenvalues at $(-\sigma_{1}^{*},\cdots,-\sigma_{v}^{*})$.
It is clear from Theorem 1 that the RADI-based approximation of $P_{\text{LQG}}$ is given by $P_{\text{LQG}}\approx\hat{V}Q_{v}^{-1}\hat{V}^{*}$.
Note that $Q_{\text{LQG}}$ is the dual of $P_{\text{LQG}}$. For completeness, we present the dual of Theorem 1 below, which produces the same approximation of $Q_{\text{LQG}}$ as RADI when the shifts are chosen as $(-\mu_{1},\cdots,-\mu_{w})$.
Theorem 2.
@@
The proof is similar to that of Theorem 1 and hence omitted for brevity. ∎
Again, it is clear from Theorem 2 that the RADI-based approximation of $Q_{\text{LQG}}$ is given by $Q_{\text{LQG}}\approx\hat{W}P_{w}^{-1}\hat{W}^{*}$.
Similar to NI-ADI-BT [61], the ADI-based non-intrusive approximation of LQGBT (NI-ADI-LQGBT) can be directly obtained using the results of Theorems 1 and 2. Specifically, by computing Cholesky-like factorizations $Q_{v}^{-1}=\hat{L}_{p}\hat{L}_{p}^{*}$ and $P_{w}^{-1}=\hat{L}_{q}\hat{L}_{q}^{*}$, and defining the low-rank factors $\tilde{L}_{p}=\hat{V}\hat{L}_{p}$ and $\tilde{L}_{q}=\hat{W}\hat{L}_{q}$, the approximations admit the factorized representations $P_{\mathrm{LQG}}\approx\tilde{L}_{p}\tilde{L}_{p}^{*}$ and $Q_{\mathrm{LQG}}\approx\tilde{L}_{q}\tilde{L}_{q}^{*}$. Substituting the exact Gramian factors $L_{p},L_{q}$ in BSA with these approximations $\tilde{L}_{p},\tilde{L}_{q}$ yields NI-ADI-LQGBT, which can be implemented using transfer function samples in the right half of the $s$-plane.
@@
The ADI-based non-intrusive approximation of $\mathcal{H}_{\infty}$BT (NI-ADI-$\mathcal{H}_{\infty}$BT) is similar to NI-ADI-LQGBT with the only difference being in the computation of $Q_{v}$ and $P_{w}$. In NI-ADI-$\mathcal{H}_{\infty}$BT, $Q_{v}$ and $P_{w}$ are computed by solving the following Lyapunov equations:
@@
3.2 Non-intrusive ADI-based Approximation of PRBT